On January 1, 2018, Pet Friendly Stores adopted the retail inventory method. Inventory transactions at both cost and retail, and cost indexes for 2018 and 2019 are as follows: 2018 2019 Retail Retail Cost Cost Beginning inventory Purchases Purchase returns Freight-in Net markups Net markdowns Net sales to customers Sales to employees (net of 25% discount) Normal spoilage Price Index: $162,500 250,000 800,000 1,084,000 $680,000 $1,063,000 4,300 7,000 12,500 2,000 2,000 12,150 6,900 4,750 950,000 22,500 4,200 11,800 8,000 722,000 22,500 6,900 1.00 1.25 1.25 January 1, 2018 December 31, 2018 December 31, 2019 Required 1. Estimate the 2018 and 2019 ending inventory and cost of goods sold using the dollar-value LIFO retail method 2. Estimate the 2018 ending inventory and cost of goods sold using the average cost retail method 3. Estimate the 2018 ending inventory and cost of goods sold using the conventional retail method
Complete this question by entering your answers in the tabs below. Req 1 Req 2 and 3 Estimate the 2018 and 2019 ending inventory and cost of goods sold using the dollar-value LIFO retail method. (Round final answers to the nearest whole dollar.) 2018 2019 Estimated ending inventory at retail Estimated ending inventory at cost Estimated cost of goods sold
Complete this question by entering your answers in the tabs below. Req 1Req 2 and 3 Estimate the 2018 ending inventory and cost of goods sold using the average cost retail method and conventional retail method. (Round your cost-to-retail percentage calculation to 2 decimal places and final answers to the nearest whole dollar.) Average Cost Conventional Retail Method Retail Method Estimated ending inventory at retail Estimated ending inventory at cost Estimated cost of goods sold Req 1 Req 2 and 3
Answer #1
Requirement 1
Employees discount must be deducted in retail column.
2018
=22500/0.75 = 30000-22500=7500 = Employee Discounts
Cost Retail
Beginning Inventory         1,62,500.00         2,50,000.00
Plus: Purchases         8,00,000.00       10,84,000.00
Freight in            12,500.00
Less: Purchases Returns             -7,000.00           -12,150.00
Plus : Net Markups              6,900.00
Less: Net Markdowns             -4,750.00
Goods available for sale (excluding beginning Inventory)         8,05,500.00       10,74,000.00
Goods available for sale (including beginning Inventory)         9,68,000.00       13,24,000.00
Base layer cost to retail percentage : =162500/250000 = 65%
2018 Layer cost to retail percentage =805500/1074000 = 75%
Less:
Normal spoilage             -4,200.00
Net sales (950000+22500)       -9,72,500.00
Employee Discounts             -7,500.00
Estimated Ending inventory at retail         3,39,800.00
Estimated ending inventory at cost         1,82,975.00
Estimated cost of goods sold         7,85,025.00
Estimated Ending inventory at cost
Ending inventory at year end retail prices Ending inventory at base year prices Inventory layers at base year retail prices Inventory layers converted to cost
                                                     3,39,800.00 =339800/1.25 = $271840         2,50,000.00 250000 x 1.00 x 65%
$162500
           21,840.00 21840 x 1.25 x 75%
=20475
Total Ending Inventory at dollar value LIFO retail cost         1,82,975.00
Employees discount must be deducted in retail column.
2019
=22500/0.75 = 30000-22500=7500 = Employee Discounts
Cost Retail
Beginning Inventory         1,82,975.00         3,39,800.00
Plus: Purchases         6,80,000.00       10,63,000.00
Freight in               2,000.00
Less: Purchases Returns             -2,000.00             -4,300.00
Plus : Net Markups            11,800.00
Less: Net Markdowns             -8,000.00
Goods available for sale (excluding beginning Inventory)         6,80,000.00       10,62,500.00
Goods available for sale (including beginning Inventory)         8,62,975.00       14,02,300.00
Base layer cost to retail percentage : =182975/339800 = 53.85%
2018 Layer cost to retail percentage =680000/1062500 = 64%
Less:
Normal spoilage              6,900.00
Net sales (722000+22500)       -7,55,400.00
Employee Discounts             -7,500.00
Estimated Ending inventory at retail         6,46,300.00
Estimated ending inventory at cost 324767
Estimated cost of goods sold 538208
Estimated Ending inventory at cost
Ending inventory at year end retail prices Ending inventory at base year prices Inventory layers at base year retail prices Inventory layers converted to cost
                                                     6,46,300.00 =646300/1.25
= 517040
        3,39,800.00 339800 x 1.00 x 53.85%
$182975
        1,77,240.00 177240 x 1.25 x 64%
$141792
Total Ending Inventory at dollar value LIFO retail cost         3,24,767.00
Requirement 2  
Employees discount must be deducted in retail column.
                                                          2,018
=22500/0.75 = 30000-22500=7500 = Employee Discounts  
Cost Retail
Beginning Inventory         1,82,975.00         3,39,800.00
Plus: Purchases           8,00,000.00       10,84,000.00
Freight in              12,500.00
Less: Purchases Returns               -7,000.00           -12,150.00
Plus : Net Markups              6,900.00
Less: Net Markdowns             -4,750.00
Goods available for sale (including beginning Inventory)         9,88,475.00       14,13,800.00
Cost to retail percentage   =988475/1413800
= 69.92%
Less:  
Normal spoilage             -4,200.00
Net sales (950000+22500)       -9,72,500.00
Employee Discounts             -7,500.00
Estimated Ending inventory at retail        4,29,600.00
Estimated ending inventory at cost(429600*69.92%)         3,00,359.92
Estimated cost of goods sold         6,88,115.08
Requirement 3
Employees discount must be deducted in retail column.
2018
=22500/0.75 = 30000-22500=7500 = Employee Discounts  
Cost Retail
Beginning Inventory         1,82,975.00         3,39,800.00
Plus: Purchases           8,00,000.00       10,84,000.00
Freight in              12,500.00
Less: Purchases Returns               -7,000.00           -12,150.00
Plus : Net Markups              6,900.00
        9,88,475.00       14,18,550.00
Cost to retail percentage   =988475/1418550
= 69.68%
Less: Markdowns               -4,750.00
Goods available for sale (including beginning Inventory)       14,13,800.00
Less:  
Normal spoilage             -4,200.00
Net sales (950000+22500)       -9,72,500.00
Employee Discounts             -7,500.00
Estimated Ending inventory at retail        4,29,600.00
Estimated ending inventory at cost(429600*69.68%)         2,99,354.17
Estimated cost of goods sold         6,89,120.83